solve the system
7x+3y=-2
-7x-7y=14

Answer by checkley71(8403)

7X+3Y=-2
-7X-7Y=14 SDDING THESE TWO EQUATIONS WE GET
-------------------
-4Y=12
Y=12/-4
Y=-3
7X+3*-3=-2
7X-9=-2
7X=7
X=7/7
X=1
PROOF
-7*1-7*-3=14
-7+21=14
14=14